OpenStudy (anonymous):

Find the equation of the line passing through (2,7) and parallel to the line passing through (1,1) and (3,-1)

OpenStudy (anonymous):

Two parallel lines have the same slope. If the line through (2, 7) is to be parallel to the line passing through (1,1) and (3, -1), then it must have the same slope. Find the slope of the line through (1, 1) and (3, -1) --> m = (-1 - 1)/(3 - 1) = -2/2 = -1. Now that you have m = -1, you can use that with the point (2, 7) and write the equation of the line in point slope form y - y1 = m(x - x1) --> y - 7 = -1(x - 2)